OpenStudy (anonymous):

What is the solution to 2x^2+x+2=0?

OpenStudy (anonymous):

use quadratic formula

OpenStudy (mathmale):

Which methods of finding solutions / roots of quadratic equations are you familiar with? Quadratic formula? Factoring? Graphing? More? Choose the one most familiar to you and apply it to the given quadratic equation. Please share your work.

OpenStudy (anonymous):

Can you take me through the steps?

OpenStudy (mathmale):

Please choose a method first. Again, pick a method (or methods) that is most familiar to you.

OpenStudy (anonymous):

I'm not really familiar with any of it. But if i had to choose one it would be, Quadratic formula, or factoring.

jagr2713 (jagr2713):

can this be factored?

OpenStudy (anonymous):

I don't know. I think so.

jagr2713 (jagr2713):

Well as you know we have to find a number that can be added to get x. But we have to but its + so i cannot be factored;. So we do have to use the quadratic formula. what are the a b c in this equation

OpenStudy (anonymous):

A is 2x^2. B is x C is 2

OpenStudy (anonymous):

?

jagr2713 (jagr2713):

Ok so we use this formula\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

jagr2713 (jagr2713):

Can you plug them in?

OpenStudy (mathmale):

We need a little housekeeping here: A is 2x^2. B is x C is 2 is not correct. a, b and c do not include the variable, x. In 2x^2 + x + 2, a is the coefficient of x^2, and is 2. b is the coeff. of x^1, and is 1. c is the constant on the far right, and is 2. In summary, a=2, b=1, c=2. Proceed, please.

OpenStudy (anonymous):

I have no idea. I'm sorry, i just don't understand. Maybe you can help me understand where to put them, so I'll know what to do from now on.

jagr2713 (jagr2713):

Well read what @mathmale just said.

OpenStudy (anonymous):

Thanks mathmale!

OpenStudy (mathmale):

Jagr has the quadratic formula correct: \[y=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (mathmale):

He will show you what to do with your a=2, b=1 and c=2.

OpenStudy (anonymous):

So i plug in the number with the letter?

jagr2713 (jagr2713):

Yes, b is x correct? so just plu in -1 for one etc etc

jagr2713 (jagr2713):

Yes @DarcyRose

OpenStudy (anonymous):

How can i get the formula so i can add the numbers?

jagr2713 (jagr2713):

We sent you the formula, here it is again\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

Like how do i write it on my computer?

OpenStudy (mathmale):

"b is x correct" is NOT correct. Please note that b is a constant, and that b = 1.

jagr2713 (jagr2713):

Oh, go to the equations button or you can just type or draw it :D

OpenStudy (mathmale):

Darcy: I'd much prefer y ou write this out on paper, or using the Draw utility (below):

jagr2713 (jagr2713):

@mjmahmood please let the user figure it out except of doing the work for them -_

OpenStudy (anonymous):

I dont even know what he wrote means lol

OpenStudy (mathmale):